CAT Chat Minutes

October 25, 2004

 

Information and Follow-up
There was an auto-page that was sent last week that announced the APS had lost power.   Floor coordinators are currently checking with CATS to see if they wish to continue receiving messages.

There was  a question if messages can be more instructive.   One user suggested that we give a minimum time to repair which would allow a user to leave and return when the problem has been corrected.   Greg Banks has stated that the MCR states the minimum time to recover at the present time.

Regarding Long Range Operations schedules that are posted to the web can a link be fixed to point to the current run on the current schedule? The current schedule is actually linked to Upcoming Schedules.

Yes, we will check the links and make any corrections necessary to show the correct current schedule.

Can a policy be made that would allow the Argonne Guest House to keep open a certain number of rooms for users?   A user called to reserve a room and was told that up to 200 rooms were currently being blocked in reserve for a conference.

We will address this request at the next OPS Directorate meeting.

General Information
Machine Stats
            Delivered beam:   417 hrs.
            Percentage of scheduled time:   98.87%
            Downtime during period:   4.77 hrs.
            Mean time between faults:   208 hrs.

Faults:

#   2   10/12 08:00   To   10/13 05:17    21.28  

 22BM PSS Fault       [SI]

  1.86 Failed processor. SI-1hr.14min, Other-37min.

#   3   10/13 07:08   To   10/17 23:05   111.95  

 Power Bump        [ ComEd]

  2.86 ComEd transformer failure, recovery, refill.

#   4   10/18 01:56   To   10/18 08:09     6.22  

  Int Dump: End of Period  

  0.05  

  0.00

#   5   10/20 08:00   To   10/25 08:04   120.07  

 Fill Ongoing             

  0.01

Beamtime Downtime Summary by APS Groups, 10/05-22/2004:

DIN

Date

Beamline

Status

Duration

Group

Description

Resolution

453

10/05/04

34-ID

Closed

3 hr 21 min

ASD-SI

Recurring Minor Fault #387 FE_PRESS_TO_0

ASD-SI installed diode across relay coil. Fault indicating insufficient air pressure in the pneumatic manifold stopped occurring. The system was tested for proper operations by ASD-SI.

456

10/08/04

31-ID

Closed

47 min

ASD-SI

BP reported that station D door close button for was not working; could not complete a search-and-secure.

PSS technician filed Work Request #16145 for adjustment of door close button hardware. Button adjusted and tested for proper operation.

460

10/12/04

24-ID

Open

7 day 0 hr 0 min

ASD-VAC

Possible pneumatic problem with PS1. FEEPS/PSS has been checked and appears to be sending the signals requesting that the shutter open. (per P. McNamara)

Replacement of PS1 solenoid.

461

10/12/04

13-ID

Closed

2 hr 12 min

ASD-SI

When users attempt to open PS2, FEEPS trips and PS1 closes. No vacuum valves close, and no PSS faults occur. (Problem occurred at 16:23 on 10/11/04 after the start of Machine Studies.)

ASD-SI fixed loose connection within PSS FERDP , that was not passing PS2 open signal to FE-EPS.

463

10/13/04

22-BM

Closed

2 day 7 hr 20 min

ASD-SI

Stored beam was lost due to Chain B fault at 22-BM. Indications on Chain B screen showed "Chain A Link Down". The beamline was taken globally offline and SI personnel called. (Initial event recorded in MCR Downtime Record #104843.)

ASD-SI replaced Chain A Processor Shelf & performed full revalidation of beamline.

464

10/14/04

01-BM

Closed

27 min

APS-PA

User cannot access Experiment Controls (computers) while PFS, on boom lifts, are in the immediate area cleaning the Visitors' Gallery windows for VIP Visit. This downtime entry is requested by CAT.

PFS completed task of cleaning windows and have left the area. Access to 1-BM control area has been allowed.

465

10/16/04

14-ID

Closed

4 hr 8 min

ASD-SI

Minor fault 442: B_DR1_CLSE_FAIL cannot be reset. PSS on-call T. Dawson is on his way in. WR #16195 submitted.

ASD-SI adjusted 14-ID-B station door. During troubleshooting, observed slow operation of Station B shutter. Time was spent ensuring that shutter functioned properly. Beamline operations restored upon completion of work.

Safety:   Beginning this Friday, training will take place for operators.   Greg Markovich, ASD, is organizing a NFPA 70E Electrical Training.   The initial classes will be scheduled; following requests should be made through the TMS representatives.